Please help, thank you!!
Randomly selected deaths of motorcycle riders in a region of the northern hemisphere are summarized in the accompanying table. Use a 0.05 significance level to test the claim that such fatalities occur with equal frequency in the different months. How might the results be explained? Month Jan. Feb. March April May June Number 9 11 13 18 19 26 Month July Aug. Sept. Oct. Nov. Dec. Number 22 29 26 15 11 6 Determine the null and alternative hypotheses. Ho: Fatalities occur with the same frequency in the different months of the year H: At least one month has a different frequency of fatalities than the other months. Calculate the test statistic, 2 (Round to three decimal places as needed.) Calculate the P-value. P.value (Round to three decimal places as needed.) What is the conclusion for this hypothesis test? Because the P-value is 5 6t 6 62 6 64 the significance level, Ho. There suficient evidence to warrant rejection of the claim that motorcycle fatalities occur with equal frequency in the66 diffarent monthe Click to select your answer(s) 67
In a clinical trial of a certain drug, 17 subjects experience headaches among the 223 subjects treated with the drug. Construct a 99% (Wald) confidence interval estimate for the proportion of treated subjects who experience headaches, completing parts (a) through (d) below a. Find the best point estimate of the population proportion. 0.076 (Type an integer or a decimal rounded to three decimal places as needed.) b. Identify the value of the margin of error E. 0.046 (Type an integer or a decimal rounded to three decimal places as needed.) c. Construct the confidence interval. eps Round to three decimal places as needed.) d. Write a statement that correctly interprets the confidence interval. Select the correct choice below and fill in the answer box to complete your choice. (Type an integer or a decimal. Do not round.) ○ A. There is a □% chance that the true value of the population proportion of subjects treated with the drug who experience headaches will fall between the confidence interval's lower and upper limits. B. We have 199% confidence that the true value of the sample proportion of subjects treated with the drm, whn ρχηerience headaches actuallv is hetwAen the nfidence interval's lower and inner limits Click to select your answer(s). iniversal
Answer #1

Following table shows the calculations:

O E=205/12 (O-E)^2/E
9 17.083 3.824555933
11 17.083 2.166065035
13 17.083 0.975875959
18 17.083 0.049223731
19 17.083 0.215119651
26 17.083 4.654503834
22 17.083 1.415260142
29 17.083 8.313228883
26 17.083 4.654503834
15 17.083 0.253988702
11 17.083 2.166065035
6 17.083 7.190358192
Total 205 204.996 35.87874893

Following is the test statistics:

(O-E35.879

The degree of freedom: df=n-1=12-1=11

The p-value using excel function "=CHIDIST(35.879,11)" is: 0.000

Since p-value is less than 0.05 so we reject the null hypothesis.

There is sufficient evidence....

----------------------------

Here we have x -17, n-223, /p--0.0762 Level of significance α=0.01 : Critical value: 2.576 The standard deviation is: POP σ,-

(c)

The confidence interval is 0.030 <p < 0122

(d)

Correct option is B.